# On the Robustness Tradeoff in Fine-Tuning

### v1

@@ Abstract
Fine-tuning has become the standard practice for adapting pre-trained (upstream) models to downstream tasks. However, the impact on model robustness is not well understood. In this work, we characterize the robustness-accuracy trade-off in fine-tuning. We evaluate the robustness and accuracy of fine-tuned models over 6 benchmark datasets and 7 different fine-tuning strategies. We observe a consistent trade-off between adversarial robustness and accuracy. Peripheral updates such as BitFit are more effective for simple tasks—over 75% above the average measured with area under the Pareto frontiers on CIFAR-10 and CIFAR-100. In contrast, fine-tuning information-heavy layers, such as attention layers via Compacter, achieves a better Pareto frontier on more complex tasks—57.5% and 34.6% above the average on Caltech-256 and CUB-200, respectively. Lastly, we observe that robustness of fine-tuning against out-of-distribution data closely tracks accuracy. These insights emphasize the need for robustness-aware fine-tuning to ensure reliable real-world deployments.
@@ 1 Introduction
Pre-training and fine-tuning can efficiently transfer knowledge from upstream data to downstream tasks [2, 43]. Models can be fine-tuned in various ways [26]—full fine-tuning, linear probing (training the classification head), and parameter-efficient fine-tuning (PEFT) [16, 29, 5, 10, 13, 15, 27, 28, 33, 48]. Specifically, PEFT strategies selectively insert/update parameters to achieve high accuracy on a targeted task with significantly reduced computation and storage costs.
@@
Prior studies on adversarial robustness [39, 35, 11, 3, 7] focus on attacking models that are trained from scratch. Models learn to use high-dimensional features during training. Since test data have similar features as training data, models are able to achieve high accuracy at evaluation. However, attackers can generate adversarial examples to largely degrade accuracy by perturbing those features, which we refer to as highly predictive, non-robust features [19, 44] in this paper. In comparison, while attacks are targeted on downstream phenomenon [4, 21, 6, 17, 38], two data phenomena are involved in pre-training and fine-tuning—upstream and downstream data, respectively. Here, a key question arises: how does adversarial robustness vary as the model is fine-tuned? We hypothesize that features learned from pre-trained data are more robust against downstream attacks. As models fit to downstream phenomenon, they learn non-robust features to gain accuracy while sacrificing adversarial robustness. Additionally, the fundamental trade-off between accuracy and adversarial robustness established in previous studies [44, 30, 50] may still exist but is potentially shifted.
This work is the first to investigate deeply how robustness is impacted by fine-tuning. Here we focus on three questions: (1) does the adversarial robustness-accuracy trade-off exist during fine-tuning? (2) how sensitive is this trade-off to different fine-tuning strategies and downstream data distributions? and (3) do findings on adversarial robustness generalize to OOD robustness? We begin by formally exploring the potential interaction between robustness and fine-tuning (i.e., the impact of mechanisms and data phenomena). Thereafter, we evaluate trade-offs empirically by fine-tuning a test suite of models using a range of diverse fine-tuning strategies (see Figure 1). The experiments are performed over 231 models, 7 fine-tuning methods, and 6 benchmark datasets (5 for adversarial robustness and 1 with 6 domains for OOD robustness), resulting in approximately 2,100 adversarial and 2,000 OOD robustness assessments.
@@
The evaluation finds: (1) a consistent adversarial robustness-accuracy trade-off early (within the first 3 epochs) in fine-tuning across all methods—initially, robustness improves together with accuracy, then it peaks and declines as fine-tuning continues; (2) the Pareto frontiers (i.e., optimal trade-off curves) are sensitive to fine-tuning methods and downstream task complexity—fine-tuning methods modifying intermediate information-intense layers, such as attention layers (e.g., Compacter [36]), achieve better balances than those updating excessively (e.g., full fine-tuning) or only peripheral layers (e.g., linear probing, BitFit [48]); and (3) OOD robustness does not exhibit a trade-off with accuracy but remains relatively stable and closely tracks accuracy, suggesting that different underlying mechanisms drive robustness in security and safety contexts. These findings deepen our understanding for designing robustness-aware fine-tuning strategies facing different tasks and risks.
@@ 2.1 Fine-tuning Strategies
ViT backbone. The transformer architecture [45] has become the state-of-the-art across many fields [9, 8, 2]. It typically serves as the backbone structure in the pre-training and fine-tuning paradigm [43], where general knowledge of pre-trained models is transferred to solve specific tasks through fine-tuning on (often small) downstream datasets. In this study, we focus on vision transformers (ViT), thus on images. Here, an input image is divided into fixed size patches, flattened, and projected into embeddings. Then attention scores are calculated to determine the relationship between patches to capture global dependencies. The outputs are then passed through a feedforward network (FFN), in-between two layer normalizations (LN) for stability.
@@
where $W_{0}$ represents the frozen pre-trained weights, and $\Delta W$ denotes the task-specific learnable parameters introduced by a given PEFT method. The input $x$ is processed through the model, and the prediction $\hat{y}$ is computed based on all parameters. Several PEFT techniques have been adapted for ViTs from the original transformer models in NLP [1]. LoRA [16] reduces computational cost by factorizing weight updates into low-rank matrices within attention layers, effectively capturing task-specific adaptations. BitFit [48] takes a more selective approach by updating only bias terms, leaving all other weights unchanged. Adapter-based methods [15] introduce small trainable modules between transformer layers to inject task-specific information without modifying the backbone. Compacter [36] further improves efficiency by using Kronecker-based parameterization within adapters, reducing the number of additional parameters needed. Finally, (IA)3 [33] fine-tunes the model by learning per-layer multiplicative reweighting factors, modifying activations without directly changing pre-trained weights. A detailed visualization on how they are applied to a ViT block is shown in Figure 2.
@@ 2.2 Model Robustness
Model robustness is crucial for evaluating the reliability of machine learning models under both adversarial manipulations and natural distribution shifts. Adversarial robustness [39, 11, 35, 3] focuses on model’s ability to defend against adversarial examples—carefully crafted perturbations that are imperceptible to humans but lead to incorrect model predictions. A widely used benchmark for security evaluation, projected gradient descent (PGD), iteratively modifies inputs to maximize model loss $L$ while constraining the perturbed image $x+\delta$ within a predefined norm-ball $\mathcal{B}$ of radius $\epsilon$ centered at the original input $x$ as shown in Equation 2.
@@ 3.2.1 Decomposition of PEFTs
To investigate the impact of fine-tuning on robustness, we select seven state-of-the-art fine-tuning strategies for our decomposition, including five PEFT methods that span all three main classes [31]: addition-based (i.e., inserting new parameters: Adapter [15], Compacter [36], and IA3 [33]), reparametrization-based (i.e., decomposing into low-rank matrices: LoRA [16]), and selection-based (i.e., modifying pre-trained weights: BitFit [48]), as well as full fine-tuning and linear probing.
@@ 3.2.2 Sensitivity Analysis
Our framework, as shown in Figure 1, systematically analyzes how robustness evolves throughout the fine-tuning process. As opposed to focusing on the final model, it captures the variances of the optimal trade-off between accuracy and robustness as the model is adapted. This is particularly important for fine-tuning. Here, the model transitions from a general to a specialized state with different numbers of robust and non-robust features learned. Encountering new data phenomena (i.e., step-level) and iterative updates (i.e., epoch-level) leads to changing degrees of robustness, as well as the trade-off.
Our approach builds upon insights from overfitting studies. There, the model’s test accuracy declines after prolonged training as it memorizes dataset-specific noises rather than generalizable patterns [37]. Similarly, fine-tuned models may learn non-robust features [19, 44] from downstream datasets to improve accuracy but degrades robustness. Here, our pipeline consists of two stages—(a) fine-tuning integration and (b) continuous evaluation. We first integrate fine-tuning modules by modifying the pre-trained model structure and/or gradients computation state. Then, as the model is fine-tuned on downstream data, we use an adaptive tracking schedule to continuously evaluate model robustness and accuracy.
Intuitively, robust and non-robust features learned from upstream and downstream data at different points have varying effects on model downstream robustness and accuracy. To capture the full variances during the training process, we monitor them at key updates during fine-tuning. A major challenge here is to determine the optimal tracking frequency. While monitoring per epoch is standard, it is too coarse for classification tasks where fine-tuned models converge within a few epochs [17]. Early-stage changes of robustness are missed with sparse tracking. Instead, we track robustness at the granularity of backpropagation steps. This captures how the robustness-accuracy trade-off evolves in two ways: (a) when new downstream-specific data is introduced, revealing immediate shifts in robustness as the model encounters new data phenomena, and (b) as the model iteratively updates after seeing the entire dataset, showing longer-term trends in robustness and accuracy trade-offs.
Furthermore, to balance efficiency with tracking at this granularity, we design a schedule that strategically samples the state of the model at selected backpropagation intervals rather than at every step. Specifically, we increase tracking frequency (i.e., every 50-200 steps) during early training, when models rapidly adapt to new data, and decrease it (i.e., every 6,000 steps) in later stages when performance stabilizes. This adaptive tracking approach ensures that we capture critical transitions in robustness while minimizing unnecessary overhead.
@@ 1. RQ1:
    Does the well-established trade-off between accuracy and adversarial robustness persist in fine-tuning?
@@ 3. RQ3:
4.1 Experimental Setup
To facilitate our experiments, we use ViT-Base model, pretrained on ImageNet-21k [42] (14 million images, 21,843 classes) at resolution 224x224 from HuggingFace [18], and AdapterHub library [1] v1.0.0 to integrate fine-tuning modules. All experiments are performed across 12 A100 GPUs with 40 GB of VRAM and CUDA version 11.7 or greater.
We fine-tune models on six different representative datasets with varying complexity (i.e., class separation, similarity with upstream data): CIFAR10 [41], CIFAR100 [24], CalTech256 [12], CUB200 [46], StanfordDogs [22], and DomainNet [40]. Detailed information about each dataset can be found in Section A.1. Configurations of the fine-tuning modules are adjusted based on common practice in solving CV tasks [14, 17] as shown in Table 1. Grid search is used to find optimal training hyperparameters. The specifics can be found in Section A.3.
@@ 4.2 Accuracy and Robustness Trade-off
The distinct trends of accuracy and robustness here highlights that models learn features for different purposes at different stages. Early fine-tuning steps improve both robustness and accuracy by leveraging pre-trained representations. We attribute this to the model’s ability to effectively adapt the randomly initialized trainable parameters to downstream tasks. Then, as the model is increasingly fitted to the downstream data, it begins to exploit predictive but non-robust features to gain accuracy while sacrificing robustness. The findings strongly confirm the persistence of the trade-off in fine-tuning and imply that the learning of robust features with defense mechanisms may lead to accuracy degradation.
@@ 4.3 Pareto Frontiers in the Trade-off Space
We attribute this variation to task complexity. CIFAR-10 consists of $10$ classes that fully overlap with ImageNet21k, which the model is already familiar with. This enables a smoother adaptation. In contrast, CUB-200 requires distinguishing 200 bird species, all falling under the broad “Bird” category of the upstream data. The complexity with smaller inter-class separation forces the model to learn finer details, increasing its dependence on fragile, non-robust features. In summary, greater downstream task complexity with less similarity with the upstream phenomena leads to steeper Pareto frontiers.
@@
Across fine-tuning strategies. Furthermore, to quantify the quality of the trade-off across fine-tuning strategies, we compute a simple scalar—area under the Pareto frontiers [32]—as our metric here. Given that Pareto frontiers capture the optimal balances that are achievable by each fine-tuning strategy, we aggregate all “suboptimal” points in that trade-off space by extending the two end points of the Pareto frontier and integrating the enclosed area, as shown in Figure 4. We refer this metric as AUC in the paper. Here, a larger value indicates a better balance between robustness and accuracy.
@@
BitFit achieves the highest AUC for CIFAR10 (75% above the average) and CIFAR100 ($81.5\%$), while Compacter outperforms others on Caltech256 by 57.5%, Stanford Dogs by 24%, and CUB200 by 34.6%. This suggests that BitFit excels on simpler datasets, where modifying only bias terms efficiently adapts pre-trained knowledge while retaining robustness. In contrast, Compacter is better for complex tasks, where its low-rank reparameterization in intermediate layers balances adaptation and robustness inheritance better. Notably, linear probing (LP) and full fine-tuning (Full FT) underperform across all datasets, with the lowest AUCs for CIFAR100 (47.5% and 20% lower than the average, respectively) and CUB200 (28.2% and 19.2%, respectively). As shown in Figure 5, linear probing and full fine-tuning result in shorter in-between Pareto frontiers, failing to achieve neither high robustness nor high accuracy compared to others. Fine-tuning all parameters destabilizes robustness, while freezing the entire network except for the classification layer limits adaptation, both resulting in suboptimal trade-offs.
@@
The observed trends align well with our preliminary results (Section 3.1). Here, the fine-tuning methods, corresponding to $k$, and downstream data phenomena, corresponding to $d$, lead to different degrees of model robustness. Furthermore, we attribute the different results of different fine-tuning methods to where (i.e., information location) and how (i.e., underlying mechanisms) the models are fine-tuned, as described in Section 3.2.1. For example, BitFit and linear probing are applied at the periphery of the model, adjusting minimal information, while LoRA, Adapter, and Compacter update deeper layers (e.g., attention mechanisms), introducing information-intense updates. Additionally, full fine-tuning adapts all parameters, leading to a more suboptimal trade-off. Here, methods that adapt intermediate representations maintain better robustness while improving accuracy, whereas peripheral or excessive updates largely degrade the balance between robustness and accuracy.

### v2

@@ Abstract
Fine-tuning has become the standard practice for adapting pre-trained models to downstream tasks. However, the impact on model robustness is not well understood. In this work, we characterize the robustness-accuracy trade-off in fine-tuning. We evaluate the robustness and accuracy of fine-tuned models over 6 benchmark datasets and 7 different fine-tuning strategies. We observe a consistent trade-off between adversarial robustness and accuracy. Peripheral updates such as BitFit are more effective for simple tasks—over 75% above the average measured by the area under the Pareto frontiers on CIFAR-10 and CIFAR-100. In contrast, fine-tuning information-heavy layers, such as attention layers via Compacter, achieves a better Pareto frontier on more complex tasks—57.5% and 34.6% above the average on Caltech-256 and CUB-200, respectively. Lastly, we observe that the robustness of fine-tuning against out-of-distribution data closely tracks accuracy. These insights emphasize the need for robustness-aware fine-tuning to ensure reliable real-world deployments.
@@ 1 Introduction
Pre-training and fine-tuning can efficiently transfer knowledge from upstream data to downstream tasks [2, 44]. Models can be fine-tuned in various ways [27]—full fine-tuning, linear probing (training the classification head), and parameter-efficient fine-tuning (PEFT) [17, 30, 5, 10, 14, 16, 28, 29, 34, 49]. Specifically, PEFT strategies selectively update parameters to achieve high accuracy on target tasks with significantly reduced computation and storage costs.
@@
Prior studies on adversarial robustness [40, 36, 11, 3, 7] focus on attacking models that are trained from scratch. Models learn to use high-dimensional features during training. Since test data have similar features to training data, models are able to achieve high accuracy at evaluation. However, attackers can generate adversarial examples to largely degrade accuracy by perturbing those features, which we refer to as highly predictive, non-robust features [20, 45] in this paper. In comparison, while attacks are targeted at downstream phenomena [4, 22, 6, 18, 39], two data phenomena are involved in pre-training and fine-tuning—upstream and downstream data, respectively. Here, a key question arises: how does adversarial robustness vary as the model is fine-tuned? We hypothesize that features learned from pre-trained data are more robust against downstream attacks. As models fit to downstream phenomena, they learn non-robust features to gain accuracy while sacrificing adversarial robustness. Additionally, the fundamental trade-off between accuracy and adversarial robustness [45, 31, 51] may still exist but is potentially shifted.
This work is the first to investigate deeply how robustness is impacted by fine-tuning. Here, we focus on three questions: (1) does the adversarial robustness-accuracy trade-off exist during fine-tuning? (2) how sensitive is this trade-off to different fine-tuning strategies and downstream data distributions? and (3) do findings on adversarial robustness generalize to OOD robustness? We begin by formally exploring the potential interaction between robustness and fine-tuning (i.e., the impact of mechanisms and data phenomena). Thereafter, we evaluate trade-offs empirically by fine-tuning a test suite of models using a range of fine-tuning strategies (see Figure 1). The experiments are performed over 231 models, 7 fine-tuning methods, and 6 benchmark datasets (5 for adversarial robustness and 1 with 6 domains for OOD robustness), resulting in approximately 2,100 adversarial and 2,000 OOD robustness assessments.
@@
The evaluation finds: (1) a consistent adversarial robustness-accuracy trade-off early (within the first 3 epochs) in fine-tuning across all methods—initially, robustness improves together with accuracy, then it peaks and declines as fine-tuning continues; (2) the Pareto frontiers (i.e., optimal trade-off curves) are sensitive to fine-tuning methods and downstream task complexity—fine-tuning methods modifying intermediate information-intense layers, such as attention layers (e.g., Compacter [37]), achieve better balances than those updating excessively (e.g., full fine-tuning) or only peripheral layers (e.g., linear probing, BitFit [49]); and (3) OOD robustness does not exhibit a trade-off with accuracy but remains relatively stable and closely tracks accuracy, suggesting that different underlying mechanisms drive robustness in security and safety contexts. These findings deepen our understanding of designing robustness-aware fine-tuning strategies facing different tasks and risks.
@@ 2.1 Fine-tuning Strategies
ViT backbone. The transformer architecture [46] has become the state-of-the-art across many fields [9, 8, 2]. It typically serves as the backbone structure in the pre-training and fine-tuning paradigm [44], where general knowledge of pre-trained models is transferred to solve specific tasks through fine-tuning on (often small) downstream datasets. In this study, we focus on vision transformers (ViT), thus on images. Here, an input image is divided into fixed-size patches, flattened, and projected into embeddings. Then attention scores are calculated to determine the relationship between patches to capture global dependencies. The outputs are then passed through a feedforward network (FFN), in-between two layer normalizations (LN) for stability.
@@
where $W_{0}$ represents the frozen pre-trained weights, and $\Delta W$ denotes the task-specific learnable parameters introduced by a given PEFT method. The input $x$ is processed through the model, and the prediction $\hat{y}$ is computed based on all parameters. Several PEFT techniques have been adapted for ViTs from the original transformer models in NLP [1]. LoRA [17] reduces computational cost by factorizing weight updates into low-rank matrices within attention layers, effectively capturing task-specific adaptations. BitFit [49] takes a more selective approach by updating only bias terms, leaving all other weights unchanged. Adapter-based methods [16] introduce small trainable modules between transformer layers to inject task-specific information without modifying the backbone. Compacter [37] further improves efficiency by using Kronecker-based parameterization within adapters, reducing the number of additional parameters needed. Finally, (IA)3 [34] fine-tunes the model by learning per-layer multiplicative reweighting factors, modifying activations without directly changing pre-trained weights. A detailed visualization of how they are applied to a ViT block is shown in Figure 2.
@@ 2.2 Model Robustness
Model robustness is crucial for evaluating the reliability of machine learning models under both adversarial manipulations and natural distribution shifts. Adversarial robustness [40, 11, 36, 3] focuses on a model’s ability to defend against adversarial examples—carefully crafted perturbations that are imperceptible to humans but lead to incorrect model predictions. A widely used benchmark for security evaluation, projected gradient descent (PGD), iteratively modifies inputs to maximize model loss $L$ while constraining the perturbed image $x+\delta$ within a predefined norm-ball $\mathcal{B}$ of radius $\epsilon$ centered at the original input $x$ as shown in Equation 2.
@@ 3.2.1 Decomposition of PEFTs
To investigate the impact of fine-tuning on robustness, we select seven state-of-the-art fine-tuning strategies for our decomposition, including five PEFT methods that span all three main categories [32]: addition-based (i.e., inserting new parameters: Adapter [16], Compacter [37], and IA3 [34]), reparametrization-based (i.e., decomposing into low-rank matrices: LoRA [17]), and selection-based (i.e., modifying pre-trained weights: BitFit [49]), as well as full fine-tuning and linear probing.
@@ 3.2.2 Sensitivity Analysis
Our framework, as shown in Figure 1, systematically analyzes how robustness evolves throughout fine-tuning. As opposed to focusing on the final model state, it captures the dynamics of the trade-off between accuracy and robustness as the model is adapted. This is particularly important for fine-tuning. Here, the model transitions from a general to a specialized state with different numbers of robust and non-robust features learned. Encountering new data phenomena (i.e., step-level) and iterative updates (i.e., epoch-level) lead to changing degrees of the robustness-accuracy trade-off.
Our approach builds on insights from overfitting studies. There, the model’s test accuracy declines after prolonged training as it memorizes dataset-specific noises rather than generalizable patterns [38]. Similarly, fine-tuned models may learn non-robust features [20, 45] from downstream datasets to improve accuracy, but it degrades robustness. Here, our pipeline has two stages—(a) fine-tuning integration and (b) continuous evaluation. We first integrate PEFT modules by modifying the pre-trained model structure. Then, as the model is fine-tuned on downstream data, we use an adaptive tracking schedule to continuously evaluate model robustness and accuracy.
Intuitively, robust and non-robust features learned from upstream and downstream data at different points influence model robustness and accuracy on target tasks. To capture the full variances during training, we monitor them at key update steps during fine-tuning. A major challenge here is to determine the optimal tracking frequency. While monitoring per epoch is standard, it is too coarse for classification tasks where fine-tuned models converge within a few epochs [18]. Early-stage changes of robustness are missed with sparse tracking. Instead, we track robustness at the granularity of backpropagation steps. This captures how the robustness-accuracy trade-off evolves in two ways: (a) when new downstream-specific data is introduced, revealing immediate shifts in robustness as the model encounters new data phenomena, and (b) as the model iteratively updates after seeing the entire dataset, showing longer-term trends in robustness and accuracy trade-offs.
Furthermore, to balance efficiency while tracking at this granularity, we strategically sample model states at selected backpropagation intervals rather than at every step. Specifically, we increase tracking frequency (i.e., every 50-200 steps) during early training, when models rapidly adapt to new data, and decrease it (i.e., every 6,000 steps) in the later stage when performance stabilizes. This adaptive tracking approach ensures that we capture critical transitions in robustness while minimizing unnecessary overhead.
@@ 1. RQ1:
    Does the established trade-off between accuracy and adversarial robustness persist in fine-tuning?
@@ 3. RQ3:
4.1 Experimental Design
To facilitate our experiments, we use ViT-Base model, pretrained on ImageNet-21k [43] (14 million images, 21,843 classes) at resolution 224x224 from HuggingFace [19], and AdapterHub [1] v1.0.0 to integrate PEFT modules. Experiments are performed across 12 A100 GPUs with 40 GB of VRAM and CUDA version 11.7 or greater. Fine-tuning models with tracking adversarial and OOD robustness takes approximately 2,100 and 3,780 GPU hours, respectively. Differences among PEFT strategies are negligible—they all update 0.07%-3.97% of the pre-trained model. Our source code is publicly available. [^1] https://github.com/kyangl/robustness-finetuning
Since fine-tuning strategies are often used in data-limited regimes, we evaluate them on six representative datasets (10k-60k samples) with varying complexity (10 vs. 256 classes, fine- vs. course-grained classes): CIFAR10 [42], CIFAR100 [25], CalTech256 [13], CUB200 [47], StanfordDogs [23], and DomainNet [41]. Dataset details are in Section A.1. To study scalability with data quantity, we also include Places365 [52] ($\sim$1.8M samples). However, as models converge to only $\sim$50clean accuracy [52] and $\sim$5adversarial robustness, their weak overall performance makes it difficult to draw meaningful conclusions about robustness trends as discussed in Section A.5. Additionally, fine-tuning configurations follow standard CV practice [15, 18] (Table 1), with grid search used to select training hyperparameters (Section A.4). We also conduct an ablation study on learning rate and the number and location of trainable parameters, which can be found in Section A.3.
@@ 4.2 Accuracy and Robustness Trade-off
The distinct trends of accuracy and robustness here highlight that models learn features for different purposes at different stages. Early fine-tuning steps improve both robustness and accuracy by leveraging pre-trained representations. We attribute this to the model’s ability to effectively adapt the randomly initialized trainable parameters to downstream tasks. Then, as the model is increasingly fitted to the downstream data, it begins to exploit predictive but non-robust features to gain accuracy while sacrificing robustness. The findings strongly confirm the persistence of the trade-off in fine-tuning and imply that learning robust features with defense mechanisms may decrease accuracy.
@@ 4.3 Pareto Frontiers in the Trade-off Space
We attribute this variation to task complexity. CIFAR-10 consists of $10$ classes that fully overlap with the pre-trained data. This enables a smoother adaptation. In contrast, CUB-200 requires distinguishing 200 bird species, all falling under the broad “Bird" category of the upstream data. The complexity with smaller inter-class separation forces the model to learn finer details, increasing its dependence on fragile, non-robust features. In summary, greater downstream task complexity with less similarity with the upstream phenomena leads to steeper Pareto frontiers.
@@
Across fine-tuning strategies. Furthermore, to quantify the quality of the trade-off across fine-tuning strategies, we compute a simple scalar—area under the Pareto frontiers [33]—as our metric. Since Pareto frontiers capture the optimal balances that are achievable by each fine-tuning strategy, we aggregate all “suboptimal" points in that trade-off space by extending the two end points of the Pareto frontier and integrating the enclosed area, as shown in Figure 4. We refer to this metric as AUC in the paper. Here, a larger value indicates a better balance between robustness and accuracy.
@@
BitFit achieves the highest AUC for CIFAR10 (75% above the average) and CIFAR100 ($81.5\%$), while Compacter outperforms others on Caltech256 by 57.5%, Stanford Dogs by 24%, and CUB200 by 34.6%. This suggests that BitFit excels on simpler datasets, where modifying only bias terms efficiently adapts pre-trained knowledge while retaining robustness. In contrast, Compacter is better for complex tasks, where its low-rank reparameterization in intermediate layers balances adaptation and robustness inheritance better. Notably, linear probing (LP) and full fine-tuning (Full FT) underperform across all datasets, with the lowest AUCs for CIFAR100 (47.5% and 20% lower than the average, respectively) and CUB200 (28.2% and 19.2%, respectively). As shown in Figure 5, LP and Full FT result in shorter Pareto frontiers, failing to achieve either high robustness or high accuracy compared to others. Fine-tuning all parameters destabilizes robustness, while freezing the entire network except for the classification layer limits adaptation, both resulting in suboptimal trade-offs.
@@
The observed trends align well with our preliminary results (Section 3.1). Here, the fine-tuning methods, corresponding to $k$, and downstream data phenomena, corresponding to $d$, lead to different degrees of model robustness. Furthermore, we attribute the different results of different fine-tuning methods to where (i.e., information location) and how (i.e., underlying mechanisms) the models are fine-tuned, as described in Section 3.2.1. For example, BitFit and LP are applied at the periphery of the model, adjusting minimal information, while LoRA, Adapter, and Compacter update deeper layers (e.g., attention mechanisms), introducing information-intense updates. Additionally, full fine-tuning adapts all parameters, leading to a more suboptimal trade-off. Here, methods that adapt intermediate representations maintain better robustness while improving accuracy, whereas peripheral or excessive updates largely degrade the balance between robustness and accuracy.